OpenStudy (selena2345):

Evaulate the expression for x=10 2(3x-15)

OpenStudy (anonymous):

2(3(10)-15) = 2(30-15) = 60-15= 45 answer is 45

OpenStudy (selena2345):

\[x=12 y=5 (x-1)2+y

OpenStudy (selena2345):

Thank You Cece Cristy

OpenStudy (anonymous):

your welcome